2. Find the complete integral of the partial differential

equation p+q=sin x+sin y.


1
Expert's answer
2021-07-19T05:43:49-0400

The given differential equation can be written as


"p-\sin x=\sin y-q"

Let


"p-\sin x=a=\sin y-q"

We have


"p=a+\sin x, q=\sin y-a"

"dz=(a+\sin x)dx+(\sin y-a)dy"

Integrate


"z=ax-\cos x-\cos y-ay+b"

The complete integral of the given PDE is


"z=a(x-y)-(\cos x+\cos y)+b"